what is the pH of a 0.150M NH4Cl solution? Kb of NH3 = 1.76x10^-5
Answer #1

c kb of N₂ 2 1 76 x 105 Kax Kb 2 10-14 Ka of NH = 10-14 = 5.68 x 1070 6 176 X10 N H + HO hotio NHA Initial C C = D. 150M conchots NHA N H + H. C = D.150M Initial enne c ( eam convinc 7 C-n 5.68 × 10 10 212 0.150-n n² = 5.68 X 10-10 (ac >>) От 150 z 9